Circular No. 7259

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V723 CASSIOPEIAE R. J. Rudy, D. K. Lynch, and S. Mazuk, The Aerospace Corporation;

R. C. Puetter, Center for Astrophysics and Space Science, University of California at San Diego; and C. E. Woodward, University of Wyoming, report 0.8–2.5-μm spectrophotometry of V723 Cas = N Cas 1995 using the Shane 3-m telescope at Lick Observatory and the Aerospace Near Infrared Imaging Spectrograph (NIRIS) on Aug. 29.5 UT. Four years after outburst, V723 Cas remains bright and displays a remarkable high-excitation spectrum. The coronal line [Si VI] 1.9629-μm is the second strongest feature in the spectrum (after He I 1.0830-μm). Two other coronal features, [Ca VIII] 2.3202-μm and [Si VII] 2.4827-μm, are also prominent. The hydrogen Paschen and Brackett series, and very strong lines of He II, are present, as are lines of [S II] and [S III]. The presence of [Si VII] 2.4827-μm requires ionizing photons with energies in excess of 205 eV. J, H, and K magnitudes determined from the spectrum are 11.9, 12.0, 11.2, respectively.

V4633 SAGITTARII = NOVA SAGITTARII 1998 Rudy et al. also report 0.8–2.5-μm spectrophotometry of V4633 Sgr =

N Sgr 1998 (cf. IAUC 6847, 6848, etc.) using the Shane 3-m telescope (+ NIRIS) at Lick Observatory on Aug. 29.5 UT. V4633 Sgr is in the early stages of its coronal phase. [Si VI] 1.9629-μm is strong and [Ca VIII] 2.3202μm is detected. He II 1.0124-μm is ∼ 2.5 times as strong as Paδ, about half the value that it will obtain as the excitation increases. Lines of [N I] and [S III] are weak but present, indicating some regions with densities in the range 10^{6} to 10^{7} cm^{−3}. J, H, and K magnitudes determined from the spectrum are 13.3, 13.2, and 12.7, respectively.

COMET C/1999 N2 (LYNN)

Ephemeris extension to IAUC 7241 (elements MPC 35554):

1999TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Sept. 19 13^{h}14.^{m}77 +33^{o}51.^{′}4 1.943 1.315 38.^{o}5 28.^{o}4 11.1
29 13 25.00 +36 02.5 2.050 1.453 41.1 27.0 11.7
Oct. 9 13 35.64 +38 17.8 2.129 1.592 45.2 26.5 12.2
19 13 46.83 +40 45.5 2.183 1.731 50.6 26.4 12.6
29 13 58.65 +43 32.0 2.217 1.868 56.8 26.4 12.9
Nov. 8 14 11.20 +46 42.4 2.236 2.004 63.7 26.3 13.3
18 14 24.58 +50 20.7 2.246 2.138 71.0 25.9 13.6
1999 September 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green